OpenStudy (sugarapples):

pi/4 please I need help medal and I will fanned them

Nnesha (nnesha):

4 pie ?

Nnesha (nnesha):

what about pi/4 ?

OpenStudy (sugarapples):

its 

Nnesha (nnesha):

:-) got that but we have t odo with it ? any statement ?

Nnesha (nnesha):

what*

Nnesha (nnesha):

change that to degree ?

OpenStudy (sugarapples):

in my book it says rewrite each degree measure in radians and each radian measures in degree

Nnesha (nnesha):

okay so formula is \[\huge\rm r \times \frac{ 180 }{ \pi }\] r= radians bec there is a pi so pi/4 it's a radians

Nnesha (nnesha):

so replace r by pi/4 and then multiply

Nnesha (nnesha):

radians to degree \[\large\rm r \times \frac{180}{\pi}\]r=radians Degree to radians \[\large\rm \frac{\pi}{180} \times d\] d=degree